### v1

# Exploring the magnetic landscape of easily-exfoliable two-dimensional materials

Abstract
Magnetic materials often exhibit complex energy landscapes with multiple local minima, each corresponding to a self-consistent electronic structure solution. Finding the global minimum is challenging, and heuristic methods are not always guaranteed to succeed. Here, we apply a recently developed automated workflow to systematically explore the energy landscape of 194 magnetic monolayers obtained from the Materials Cloud 2D crystals database and determine their ground-state magnetic order. Our approach enables effective control and sampling of orbital occupation matrices, allowing rapid identification of local minima. We find a diverse set of self-consistent collinear metastable states, further enriched by Hubbard-corrected energy functionals, when the parameters have been computed from first principles using linear-response theory. We categorise the monolayers by their magnetic ordering and highlight promising candidates. Our results include 109 ferromagnetic, 83 antiferromagnetic, and 2 altermagnetic monolayers, along with 12 novel ferromagnetic half-metals with potential for spintronics technologies.
Magnetism in two-dimensional (2D) van der Waals (vdW) materials has garnered significant attention in recent years, 1; 2; 3; 4; 5; 6; 7; 8; 9 not only for potential technological applications but also for the large variety of phenomena that can be hosted by magnetic 2D materials. Indeed, magnetism in 2D displays very different properties depending on the interplay between thermal fluctuations, exchange interactions, and magnetic anisotropy. The thrust to investigate this richness in experiments has stimulated an intense search for novel magnetic monolayers displaying specific phenomena, from Ising anisotropy, 10 to Berezinskii-Kosterlitz-Thouless transitions, 11 and heavy-fermion excitations 12.
Although many candidates have been suggested by a long-term experience with layered magnetic materials, the continuous demand for novel 2D magnets calls for more systematic and exhaustive investigations. In this respect, theoretical approaches, particularly first-principles calculations based on density functional theory (DFT), provide a robust and high-throughput approach to predict novel 2D materials, 13; 14; 15; 16; 17; 18; 19; 20; 21; 22 including magnetic monolayers 13; 23; 24 and their properties. 25; 26
Despite many successes, DFT can sometimes struggle to accurately predict the magnetic properties of materials, owing to the localized nature of - and -orbitals typically involved in magnetism and their tendency to suffer from self-interaction errors when adopting approximate exchange-correlation functionals. A computationally inexpensive and effective strategy to overcome these limitations is achieved by complementing approximate functionals with self-interaction corrections 27 in the so-called Hubbard DFT+ approach. 28; 29; 30 The Hubbard parameter entering the DFT+ approach can be considered, somehow unsatisfactory, as a semi-empirical parameter, or can be computed self-consistently, e.g. using linear-response theory 31. In the latter case, the theory preserves a first-principles character, and efficient implementations using density-functional perturbations theory (DFPT) are suitable for high-throughput investigations. 32; 33; 34; 35; 31; 36; 30; 37; 38
Another crucial challenge for simulations is the identification of the ground state magnetic configuration of the system. Indeed, the magnetic energy landscape is typically complex, with several local minima associated with different magnetic states. The inclusion of Hubbard corrections, needed for a better description of magnetic systems, typically leads to a proliferation of local minima, making it even more challenging to identify the true ground state (global minimum).39; 40; 41; 42; 43; 44; 45; 46; 47; 48 Several approaches have been developed to tackle this issue, ranging from efficiently generating different magnetic configurations, 49 to the manipulation of the occupation matrix for - or -orbitals, 50; 51; 52; 53; 39 to computing spin waves 54 and spin spirals 55, employing cluster multiple theory, 56 or even genetic algorithms. 57
In this work, we thus explore the magnetic properties of easily exfoliable 2D materials from the Materials Cloud two-dimensional crystals database (MC2D) 14; 13; 58; 59 using our recently developed approaches to reveal novel magnetic monolayers. Starting from 3077 easily and potentially exfoliabble monolayers (2004 easily and 1073 potentially exfoliable) with up to 40 atoms per cell 14; 13, 877 systems with up to 12 atoms per cell are screened with plain DFT using an AiiDA60; 61 workflow (Chronos)13. Out of these systems, 483 monolayers are found to be non-magnetic, 166 systems were discarded because of the failure in the optimization, and 228 monolayers showed a magnetic ground state. For these 228 materials, Hubbard corrections are then included for better accuracy, with the Hubbard parameter computed self-consistently. To identify the magnetic ground state, we employ the Robust Occupation Matrix Energy Optimization (RomeoDFT) algorithm 48 based on constraining the atomic orbital occupation matrices that allows us to systematically explore the magnetic energy landscape. With this protocol, we identify 194 magnetic 2D materials for which we accurately determine the magnetic ground state and investigate electronic and magnetic properties, including a proxy for the critical temperature. Notably, we identify 12 half-metals, a class of materials with spin-dependent conductivity, where electrons in one spin channel (up or down) are metallic, while those in the opposite channel are insulating, which is especially promising for spintronics applications. The remaining 34 systems were discarded because either they were repeated, found non-magnetic, consistently failed in the calculation of the Hubbard parameter, or not enough minima were found using RomeoDFT (see Supporting Information).
Results and Discussions
To identify magnetic 2D materials and their magnetic ground state, we consider monolayers from MC2D,58; 59 a portfolio of 2D materials that can be exfoliated from experimentally known 3D parent compounds, identified using high-throughput computational exfoliation. 13; 14 We start from 877 easily exfoliable 2D materials with up to 12 atoms per unit cell. As summarized in Fig. 1, the tendency to show magnetism is first tested using an improved version of the AiiDA 60; 61 workflow, named Chronos, originally adopted in Ref. 13 to screen a smaller set of materials. For each system, the workflow first considers several ferromagnetic configurations through different starting magnetizations on the atoms and computes its optimized geometry and total energy through collinear DFT calculations with an ordinary approximate exchange-correlation functional (PBE). If any ferromagnetic state is found to have a total energy lower than a non-magnetic reference calculation, the system is considered potentially magnetic and additional magnetic configurations are investigated, including antiferromagnetic states that require up to a supercell to accommodate two magnetic atoms with opposite spins (see Methods for more details). In this way, 228 magnetic monolayers are identified with up to 12 atoms in their primitive cell, together with their potential collinear ground state.
Out of these, 174 are found to be ferromagnetic, and further broken down into 100 metals and 74 insulators. The remaining 54 materials show an AFM ground state, with 31 having a finite gap and 23 being metallic.
Although versatile and flexible, the workflow explores possible magnetic states only through the starting magnetization, that is, through initial conditions on the density of spin up and spin down states on each atom. A more robust way is to initialize from the atomic orbital occupation matrices - particularly and orbitals where the magnetization usually stems from - and explore more possible states. Introducing atomic orbitals labeled by on the -th atom, we can define an atomic occupation matrix as , where are the occupations of Bloch states and magnetization is an unbalance between the occupation of spin-up and down states . To explore more systematically the magnetic energy landscape and obtain a more reliable identification of the magnetic ground state, it is important to gain full control over the atomic occupation matrix (at least for the relevant atomic orbitals on the magnetic atoms), not just through the spin density. 50; 51; 52; 53; 39 We thus validate the magnetic ground-state identification by applying a constraint on the atomic orbital occupations and then letting the self-consistent procedure find the closest local minimum in the energy landscape. The overall global minimum corresponding to the magnetic ground state can then be found by exhaustively exploring possible atomic occupations and comparing the energies of the corresponding self-consistent solutions. To perform this task automatically and efficiently, we use the RomeoDFT workflow 48 (see Methods for more details). We first consider a subgroup of magnetic monolayers and find that without Hubbard corrections in essentially all cases the same ground state is found either by controlling the starting magnetization by spin density or the constraints on the full atomic occupation matrix (see figure S1 in Supporting Information. When considering PBE calculations, the search implemented in the Chronos workflow is thus exhaustive enough to reproduce the more sophisticated exploration performed by RomeoDFT.
Most of the identified 2D magnetic materials host electrons in localized or orbitals, for which approximate DFT functionals such as PBE tend to suffer from inaccuracies arising from self-interaction. 27; 62 The magnetic ground-state predictions might be affected by such errors and need to be validated against more accurate calculations. As alluded to before, an approach to mitigate self-interaction problems, known as DFT+, relies on complementing DFT functionals with terms inspired by the Hubbard model. 63; 64; 65; 28; 66; 67; 68; 30; 69 Such Hubbard-corrected functionals have been shown to remove prominent self-interaction errors when the Hubbard parameter is computed self-consistently from linear-response theory, 31 thus preserving the parameter-free character of DFT. Unfortunately, the inclusion of Hubbard corrections makes the energy landscape much more complex, with the appearance of many emergent local minima with similar total energies. 48 To illustrate this effect, we consider the case of monolayer CoO2 in Fig. 2, where each point represents a local minimum in the energy landscape, identified by exploring multiple occupation matrices. At the PBE level (), only a limited set of local minima is found, with an overall ferromagnetic ground state (indicated by a star). When including Hubbard corrections, with the Hubbard calculated self-consistently from the linear-response method (PBE+), the number of local minima significantly increases, making the magnetic energy landscape more complex and the possibility of falling in a local minimum different from the true ground state higher. We thus have that, while at the PBE level () the number of states is limited and the identification of the global ground-state minimum is simpler (explaining the agreement between Chronos and RomeoDFT at the PBE level (figure S1 in the Supporting information)), an approach based on the occupation matrix is mandatory when Hubbard corrections are included. In Fig. 2 we also show results for an intermediate case with eV, where the proliferation of local minima is already present, although less dramatic than at . Remarkably, the ground state, labeled by a star in each panel, is very sensitive to the value of as the system is predicted to be ferromagnetic at , antiferromagnetic at 4 eV and then ferromagnetic again (but with a different magnetization) at eV. 70 Therefore, the choice of strongly affects the ground-state properties of the system. To avoid any empiricism, we compute the Hubbard parameter for each material using an efficient formulation of the linear-response approach 31 within density-functional perturbation theory 33; 34; 35 (see Methods and Supporting Information for more details).
Equipped with self-consistent Hubbard corrections, we navigate the magnetic energy landscape of potentially 2D magnetic materials identified through the Chronos workflow at the PBE level, now using PBE+ and a control over the atomic occupation matrix using RomeoDFT. Out of 228 materials, 34 systems are discarded because either they are found to be nonmagnetic or there are convergence issues in the calculation of the Hubbard parameter or self-consistent calculations, leaving us with 194 magnetic monolayers for which we characterize their ground state. The results are summarized in Fig. 3, where we report the absolute magnetization per magnetic atom and the band gap of the ground state of all systems, together with their distribution. More details on the specific magnetic configuration, crystal structure, magnetic energy landscape, and band structure for each material are provided in the Supporting Information.
We categorize all systems into four groups, depending on whether they are ferromagnetic or antiferromagnetic, insulators or metals (we did not find ferrimagnets, and the two altermagnets are included in the antiferromagnetic group).
fig. 4 shows the number of systems in each group, with the left and right charts representing results from PBE (Chronos workflow) and PBE+ (RomeoDFT workflow) calculations, respectively. Blue and red colors indicate antiferromagnetic and ferromagnetic materials. The innermost rings show the number of systems in each category, while the second rings count metallic (hatched) or semiconducting (solid) systems. In the right chart, the outermost ring indicates the number of systems from the innermost ring that either remained in the same magnetic category (solid fill) or changed categories (dotted fill) from the PBE calculations. As expected, the inclusion of Hubbard corrections leads to an increase in the number of insulating systems. From fig. 4, we also note that using the RomeoDFT workflow with PBE+ calculations significantly expands the number of antiferromagnetic systems (blue colors): approximately 40% of these materials were identified as ferromagnetic when using the Chronos workflow with standard PBE calculations. This change can be attributed primarily to two factors: first, the limited capability to explore the energy landscape when using simply a starting magnetization; second, the possible increase in the number of magnetic states after including Hubbard corrections.
We note in passing that among the AFM identified above, we find only two altermagnetic monolayers by generalizing the protocol for 3D systems 71 to include additional symmetry constraints in 2D. 72; 73 These are RuF4 and VF4, with the same structural prototype. Both systems were already predicted to be 2D altermagnets in Ref. 72; we do not see altermagnetism in FeBr3, because its crystal structure and symmetry is different from the one studied in Ref. 72.
Due to the scarcity of experimental results for magnetic monolayers, it is challenging to validate our predictions against experiments. We thus focus on a few cases reported in the existing literature. For example, our results agree with the experimentally observed ferromagnetism in the family of chromium trihalides (CrX3; X = Cl, Br, I with CrCl3 being 2D-XY ferromagnetic) 74; 10; 11 and in CrSBr. 75
In the family of transition-metal phosphorus trisulfides, our results match the experimentally observed ferromagnetism in CrPS4 76 and the antiferromagnetic ordering in NiPS3, although suppressed by fluctuations in the monolayer limit. 77 For the case of FePS3, our findings show ferromagnetic ordering, in contrast with the antiferromagnetic ordering observed in experiments. 78 We attribute this discrepancy to the limited size of the unit cell considered in this work for FePS3, which is not able to accommodate the measured zigzag ordering. Indeed, since the primitive cell already contains two magnetic atoms, no supercell is considered, and only a Néel AFM ordering is tested. This limitation, combined with the use of collinear calculations, could influence the accuracy in predicting other antiferromagnetic systems and may lead to discrepancies with other studies, 79; 55, and we expect it to be at the origin of the relatively small number of AFM systems with respect to FM.
Among the possible conclusions that we can draw from the high-throughput investigation of the magnetic ground-state of 2D materials, we also focus on the nature of magnetic elements in each structure. Conventional magnetism usually arises from or orbitals, qualifying transition metals and lanthanides as the dominant contributors to magnetism. Our results largely confirm such conventional wisdom, as shown in fig. 5, where we report the number of systems in which a given element is magnetic. However, fig. 5 also suggests that certain halogen and nonmetal elements (especially oxygen) also play a role in magnetism in 2D compounds, putting forward their potential role in unconventional magnetism. For instance, in GaSe (space group: ) and SiN2F6, the selenium and nitrogen atoms exhibit non-negligible magnetic moments (0.2 and 2 respectively). Interestingly, in the case of CdOCl, while the orbitals of Cd are fully occupied and exhibit negligible magnetic moments, magnetism predominantly arises from the oxygen atoms, which have a magnetic moment of 0.8 .
Exploring the energy landscape gives us access not only to the ground state of the system but also to the ability to quantify the energy difference between the ground state and other metastable magnetic states. This is particularly relevant when it is possible to identify a metastable state with magnetic moments on the atoms similar to the ground state but different magnetic ordering, e.g., a ferromagnetic ground state and a metastable antiferromagnetic state with comparable absolute magnetization (or vice-versa). The energy difference between these states can then be translated into an effective exchange parameter of an underlying Heisenberg model through
where is the number of magnetic atoms per cell while and are the energies of the ferromagnetic and antiferromagnetic states, respectively. This effective exchange parameter is related to an effective nearest-neighbor exchange coupling through , where is the number of nearest neighbors. A large energy difference, , corresponds to a large exchange parameter , stabilizing the magnetic order and possibly resulting in a higher critical temperature. It is important to note that the mapping between first-principles calculations and the Heisenberg model is meaningful only when the magnetization arises from local atomic magnetic moments that have the same values (although different orientations) in the FM and AFM states, i.e., SFM = SAFM. This condition is typically met in insulators, while in general it cannot be readily applied to magnetic metals, where magnetism arises from itinerant electrons. Therefore, we consider only magnetic insulators to compute and the effective .
fig. 6 shows divided by the number of magnetic atoms per cell, , for the insulating magnetic monolayers for which it was possible to find a metastable magnetic state with similar atomic magnetic moments but different ordering. Systems are sorted with increasing effective . The bottom panel contains systems where exchange interactions are weak, making it more challenging for the magnetic ordering to compete with thermal fluctuations. On the contrary, the systems in the top panel are expected to have higher critical transition temperatures. The ranking based on the effective seems to reproduce several experimental trends for the ordering temperature. For instance, it is experimentally known that CrSBr has a higher critical temperature than the chromium trihalides, 10; 75 and within this family the iodine compound has the largest while the chlorine one the smallest, 80; 81; 82; 10; 11 in agreement with the ordering in fig. 6.
To further validate as a meaningful proxy for , we select three systems with the largest in fig. 6 and compute their transition temperatures using different approaches. We thus focus on MoCl3, GdI2, and VPS3, noting that the latter was already identified as a potentially high- magnetic monolayer based on high-throughput calculations. 83 Additionally, we include CrI3 in our analysis, as experimental results are available for comparison. In addition to the approach based on eq. 1 and relying on the energy difference between FM and AFM reference states, we also consider the so-called four-state method, 84 which calculates the exchange parameter based on the energy differences between four distinct magnetic configurations of a pair of neighboring atoms in a supercell (see Methods for more details). Remarkably, the four-state method allows the calculation of specific interatomic exchange parameters, not just an effective nearest-neighbor as in eq. 1, and we consider interactions up to the third (fourth) nearest neighbors for CrI3 and GdI2 (VPS3 and MoCl3). The exchange parameters computed within the two approaches are then denoted and in the following. From the knowledge of the exchange parameters, the transition temperature is calculated using either mean-field (MF) theory or classical Monte Carlo (MC) simulations. In the latter, we completely neglect relativistic effects. Indeed, magnetocrystalline anisotropy does not affect significantly the critical temperature for a finite-size system, and even isotropic monolayers can sustain magnetic order at finite temperature when considering realistic sample sizes. 85
Critical temperatures computed either using MF or MC, with exchange parameters estimated from both eq. 1 and the four-state method, are shown in fig. 7 for CrI3, GdI2, VPS3, and MoCl3. Trends in derived from both MC and MF methods are consistent with fig. 6, with MF calculations showing higher values, as expected from the well-known overestimation of in mean-field theory. 86
We note that, in general, different methods for computing the exchange parameters give similar values of . The only exception is MoCl3, for which the MC estimates of differ significantly between the two methods, with a notably lower value when the exchange parameters are calculated using the four-state method. This discrepancy is attributed to the dimerization of Mo atoms in MoCl3, which distorts the otherwise honeycomb lattice of Mo atoms. The energy difference between the FM and AFM is mainly arising from a strong AFM coupling between dimerized pairs, while dimers are only weakly coupled. This feature is correctly captured within the four-state method, while the effective approach of eq. 1 uniformly distributes the FM-AFM energy difference between the closest atoms. This failure also affects all MF estimates in a similar way. We thus expect that the effective reported in fig. 6 provides a good estimation of the ranking in critical temperature, provided that exchange interactions are distributed evenly between similar neighbors.
We also note that the transition temperature of the CrI3 monolayer is higher than the values reported in previous studies and experiments87; 25; 10. This discrepancy can be attributed to differences in geometry, the relatively large value for CrI3 obtained from linear-response,38, and the influence of broadening effects in Brillouin-zone integrations. For consistency, we employed the same broadening value as used in the Chronos workflow (270 meV). Since the exchange parameter is derived from total energies–which are sensitive to the choice of broadening–it is likewise influenced by the selected broadening value.
Among metallic systems, we are particularly interested in half-metallic ferromagnetic systems that are interesting for spintronics devices because they can provide fully spin-polarized currents. 88 We thus search for systems where the Fermi energy crosses the energy bands for one spin channel while the other spin channel remains gapped. 89 Twelve novel half-metals are identified, whose electronic structure is reported in fig. 8. In several cases the bandwidth is sufficiently large to expect spin-polarized transport even in the presence of disorder. Remarkably, in systems like CuO2, Cr3O8, EuOI, EuOBr2, VO3, VS2O8, YbBr3, and YbCl3 a large gap is present above the metallic spin channel, which is potentially interesting in view of possible optoelectronic control of spin currents. In all these cases a correct description of the magnetic ground state is crucial to have half-metallic properties. For instance, in the case of FeAl2S4 it is important to have the two iron atoms with different oxidation states (+1 and +2) and thus different magnetic moments of 2.2 and 3.2 . We also stress that, in general, the half-metallic state can be fragile, especially if it is protected by symmetry-induced degeneracies. 90 Further tests to include the effect of spin-orbit coupling or structural distortions would thus be needed to fully validate predictions, although the procedure outlined above is already sufficient to discard speculative half-metals such as iron dihalides FeX2 (with X = Cl, Br, I). These systems have been predicted to be half-metals, 91; 92 although experimental evidence 93; 94; 95; 96; 97suggests that they have a finite gap. We find them to be indeed insulating, although the mechanism to break the symmetry-induced degeneracy arises from an incorrect AFM configuration, as we don’t include relativistic effects or distortions.
I Conclusions
In conclusion, we investigate the magnetic ground states of easily exfoliable monolayers from the MC2D database. Starting from 3077 easily and potentially exfoliable monolayers identified in Refs. 14; 13, an initial screening is carried out on 877 easily-exfoliable systems containing up to 12 atoms per cell using high-throughput density functional theory with the conventional PBE exchange-correlation functional. Testing various magnetic configurations by changing the initial magnetization, 228 magnetic monolayers are identified as potentially magnetic systems. Subsequently, Hubbard corrections are introduced for these systems, with the Hubbard parameter determined self-consistently. The magnetic energy landscape is then thoroughly explored by applying constraints to the occupation matrices of the or orbitals, enabling the identification of the global energy minimum. Through this process, we identify 194 magnetic monolayers, comprising 109 ferromagnetic, 83 antiferromagnetic, and 2 altermagnetic systems. To estimate critical temperatures and highlight promising candidates, we employ a proxy based on the energy difference between ferromagnetic and antiferromagnetic states. This method is applied to the 137 monolayers that exhibit an electronic band gap. Among the remaining 57 metallic systems, we identify 12 novel half-metals with strong potential for spintronics applications.
Methods
In this work, all calculations are collinear and performed using the Quantum ESPRESSO (QE) distribution. 98; 99; 100 Calculations are performed within the
spin-polarized generalized-gradient approximation (GGA) for exchange-correlation functional using the Perdew-Burke-Ernzerhof (PBE) prescription. The pseudopotentials and the wavefunction kinetic-energy cutoffs are taken from a standard solid-state pseudopotentials (SSSP Efficiency) library. 101; 102; 103; 104; 105 We use the most recent versions of the SSSP library available at the time for our calculations. Specifically, version 1.1.2 is used for the PBE calculations, and version 1.3.0 is utilized for the PBE+ calculations. The -points density of and -points mesh density of have been set for single self-consistent (scf) and density functional perturbation theory (DFPT) calculations, respectively. In this work, a system (an atom) is magnetic if the absolute total magnetization per cell (of the atom) is larger than 0.1. The data used to produce some of the results of this paper are given in the materials cloud archive. 106
In this work we consider systems from MC2D, 58; 59; 14 which is created from a collection of 3077 exfoliable monolayer that were identified from prototypes of the Inorganic Crystal Structure Database (ICSD), 107; 108 the Crystallography Open Database (COD), 109 and Materials Platform for Data Science (MPDS) 110 databases. We identify and characterize magnetic monolayers in MC2D in two main steps, as shown in fig. 1: In the first step (Chronos step), using PBE calculations, potentially magnetic monolayers are identified, and in the second step (RomeoDFT step), after including Hubbard corrections and exploring the energy landscape, the magnetic ground state of these materials is found. An initial screening was performed using an automated AiiDA 60; 61 workflow (Chronos) on 877 easily exfoliable materials (i.e. materials with a binding energy per unit of area lower or equal to 30 meV/Å2 when computed with the vdW-DF2-c09 functional or lower or equal to 35 meV/Å2 when computed with the rvv10) with up to 12 atoms per unit cell, using PBE calculations, and 228 magnetic monolayers are identified by following procedure: the geometry of the system is optimized in ferromagnetic and non-magnetic orderings. If the ferromagnetic ordering shows lower energy than the non-magnetic case, the system is considered to be potentially magnetic, and other antiferromagnetic orderings (making sure the supercell size is now to accommodate two magnetic atoms) and five random configurations are created. The system is optimized starting from different starting magnetization defined by spin density ( where describes the density and I is the atom index). As stated before this simplistic approach is effective in capturing the correct magnetic ground state in the absence of Hubbard correction (see Supporting Information). However, it is important to note that there are still local minima within the energy landscape that are accessible through the manipulations of the orbital occupations.
This can be achieved using RomeoDFT, which corresponds to the second step of our overall workflow. In this approach, the system is induced towards a set of target occupation matrices using Lagrange multipliers:
with
where EU is the Hubbard corrected energy, is the Lagrange multiplier, representing the strength of the energy penalty associated with a deviation from the target occupation matrices, is the atomic site index, and are the magnetic quantum numbers associated with a specific angular momentum and is the on-site Hubbard parameters. are the generalized atomic occupation matrices which are computed by projecting the Kohn-Sham (KS) wavefunctions on atomic orbitals as: , where are the occupations of KS states. The general formulation of DFT+ is discussed in Refs. 30; 37; 35; 33. We note that in this case the magnetization is defined from the occupation matrix of spin up and down: . After imposing the constraint for a fixed number of self-consistent cycles, the constraint is released and the system is allowed to evolve freely to the closest self-consistent energy minimum. More information about RomeoDFT is provided in Ref. 48. This implementation is made in Quantum ESPRESSO and the process is automatized in Julia and can be found on GitHub. We use RomeoDFT workflow to determine the ground state at the PBE+ level for 194 magnetic monolayers, identified by Chronos with a filter of up to 12 atoms per primitive cell.
In order to calculate the parameter, we use DFPT 35; 33 as implemented in the HP code, 34; 31; 37 which is part of QE. The Hubbard parameter is computed in three iterative steps. Within each iteration the value is updated as the average values of the initial and calculated values in the previous step. 111Only for special case of CoO2, because the ground state is highly sensitive to the choice of , in order to avoid falling into the wrong states we do not update the value as the average of the previous step, instead, in each step we use the same calculated value of the from the previous step. Our results show that with this approach, the convergence threshold of for 87 of the systems is less than 0.5 eV. If the convergence of 0.5 eV has not been reached in three steps, we do another final step. After the calculation of the , the geometry and the Hubbard are fixed, and we explore the energy landscape to find the global minimum. For the systems with fully occupied orbitals we use a zero value for the because the effects of the Hubbard correction at integer occupations are negligible. 31
In the entire process the geometry is fixed to the ground state as identified by Chronos prior to optimization to avoid introducing geometric optimization only on one state. The appropriate approach involves relaxing the geometry across all states within the energy landscape, which can be computationally demanding.
To further characterize the magnetic properties of our materials, we mapped their electronic structure onto a model Heisenberg Hamiltonian:
where corresponds to the magnetic exchange coupling between the net spins in sites and . The magnitude of the spins is included in the . The sum runs over every spin such that each coupling is counted twice, i.e. and are both included. Our convention is such that a negative indicates a ferromagnetic coupling, and a positive is indicative of antiferromagnetic interaction. The transition temperature is calculated from mean-field theory through the formula:
where is the Boltzmann constant and . We calculated the exchange parameters using two approaches. First, we employed the method, which maps this total energy difference onto a nearest-neighbor-only model Hamiltonian and the results are shown in fig. 6. Here, the exchange coupling is assumed to be present between equivalent nearest neighbors () and given by , where is the energy difference between self-consistent ferromagnetic and antiferromagnetic spin configurations and the number of magnetic atoms in the unit cell. To account for interactions beyond the first nearest neighbors, we employ the four-state method. In this approach, a large supercell is considered (), and each exchange interaction is calculated by computing the total energy when the pair of spins in sites and are in four distinct collinear spin configurations while keeping all other spins ferromagnetically ordered. The four configurations are up-up, down-up, up-down, and down-down concerning the quantization axis. The exchange parameter is then calculated by: 84
Finally, to go beyond the mean-field theory and obtain a more precise estimate of the paramagnetic critical temperatures, we use the Monte Carlo Metropolis algorithm as implemented in the Vampire package 112; 113 for a few selected materials. Here, we use periodic boundaries in 2D, an equilibration time of 5.000 Monte Carlo steps, and statistical averaging also over 5.000 steps. The critical temperatures are then obtained by fitting the Monte Carlo magnetization curves to the mean-field formula , where and the critical exponent are also fitting parameters
It is important to note the approximations involved in our results. Besides assuming a magnetism driven by localized magnetic moments, the mapping of the onto the Heisenberg model considers only nearest neighbors coupling. This can be regarded as a mean-field approximation. Take the case of MoCl3, where the Mo atoms go from a honeycomb lattice arrangement, with three equidistant nearest-neighbor sites, to a triangular lattice of dimers where the distance between Mo sites in the dimer is 1.06 Å smaller than between dimers (see Supporting Information).
Anisotropy is also neglected in our calculations, and all calculations are done within the non-relativistic limit. While anisotropy could play a significant role in 2D materials, our MC simulations on a model honeycomb lattice considering only the nearest neighbors indicate that is not affected significantly by a typical magnetocrystalline anisotropy. This aligns with recent research, that shows that even for isotropic 2D systems in lab setups short-range interactions can be large enough to stabilize the magnetic order at a finite temperature due to finite effects. 85
Acknowledgements.
II acknowledgments
We thank Guoyuan Liu for fruitful discussions. We acknowledge support from the Swiss National Science Foundation (SNSF), through grant 200021-179138, and its National Centre of Competence in Research (NCCR) MARVEL (grant number 205602). MG acknowledges financial support from Ministero Italiano dell’Università e della Ricerca through the PNRR project ”Ecosystem for Sustainable Transition in Emilia-Romagna” (ECS_00000033-ECOSISTER) and through the PRIN2022 project SECSY (CUP E53D23001700006), both funded by the European Union – NextGenerationEU. Computer time was provided by the Swiss National Supercomputing Centre (CSCS) under project No. s1073. The authors declare no competing financial interests.